OpenStudy (anonymous):

A science experiment begins with a metal at -100° Celsius. The following function describes the temperature change per minute: f(x) = 89x - 100°. How will the graph of this function change if the metal is at 25° at the start of the experiment? The line will shift vertically up by 125°. The line will shift vertically up by 25°. The line will shift vertically down by 125°. The line will shift vertically down by 25°.

OpenStudy (anonymous):

@Andras

OpenStudy (anonymous):

Instead of f(x)=89x-100 it will be f(x)=89x+25 That is equivalent of shifting the line vertically up by 125. (-100 to 0 is 100 and 0 to 25 is an extra 25, thus 125)

OpenStudy (anonymous):

How come it changes to a plus? @Andras

OpenStudy (anonymous):

As that was stated. at 25° means +25°

OpenStudy (anonymous):

Oops sorry I didn't see that :( I have two more and they're pretty tough. Up for the challenge I'll give medals?

OpenStudy (anonymous):

As you wish I am getting tired of maths a bit. Been on this site today for too long

OpenStudy (anonymous):

Ok here goes.

OpenStudy (anonymous):

Given f(x) and g(x) = f(x) + k, look at the graph below and determine the value of k. k = ______

OpenStudy (anonymous):

@Andras

OpenStudy (anonymous):

5

OpenStudy (anonymous):

The answer is 5?

OpenStudy (anonymous):

Yes, I am getting lazy :) Do you see why?

OpenStudy (anonymous):

Not really :(

OpenStudy (anonymous):

What is common in the 2 lines?

OpenStudy (anonymous):

they're parallel

OpenStudy (anonymous):

Yep. Let us say that the first lines equation is as follows y=mx+a

OpenStudy (anonymous):

Than what do we know of other parallel lines with this?

OpenStudy (anonymous):

Slope is the same

OpenStudy (anonymous):

y=mx+a which of these represents the slope?

OpenStudy (anonymous):

The blue??

OpenStudy (anonymous):

from the above equation

OpenStudy (anonymous):

I'm getting lazy oops

OpenStudy (anonymous):

Ok I have one last question for @Andras

OpenStudy (anonymous):

A cell phone plan has a monthly cost that is shown in the table below. What is the correct statement regarding the average rate of change during the 40 minute time of talk? Total minutes of talk time Monthly cost of cell phone 0 $19.95 10 $20.45 20 $20.95 30 $21.45 40 $21.95 The average rate of change is $0.50, meaning that for each minute of talk time, the monthly bill increases by $0.50. The average rate of change is $0.50, meaning that for every ten minutes of talk time, the bill increases by $0.50. The average rate of change is $0.05, meaning that for each minute of talk time, the monthly bill increases by $0.05. The average rate of change is $0.05, meaning that for every ten minutes of talk time, the bill increases by $0.05.

OpenStudy (anonymous):

Anyway what I wanted you to say is m. And the line was shifted vertically up by 5 thus the answer is 5

OpenStudy (anonymous):

Ohhhh, wow. Sorry I didn't readv through.

OpenStudy (anonymous):

Am I right that the question states that for nearly 20 bucks you get 0 minutes. ROBBERY

OpenStudy (anonymous):

Yes you are.

OpenStudy (anonymous):

Average! So we only need to check the end points. for 0 min: $19.95 for 40 min: $21.95 Also we are asked for the rate of CHANGE the change is 21.95-19.95=2 but that is for 40 min For 1 min it is 2/40=0.05

OpenStudy (anonymous):

So its C or D?

OpenStudy (anonymous):

OR for every 10 minutes the change is 10*0.05=0.5

OpenStudy (anonymous):

What do you think

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

FAIL

OpenStudy (anonymous):

Aw then C

OpenStudy (anonymous):

B or C. I think they are looking for B as all the changes were given in 10 minutes periods

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Yes I am sure that it is B. As the plan does not change for every single minute

OpenStudy (anonymous):

Alright man! I'm done.